### v2

# Diffraction by a Right-Angle Penetrable Wedge: Closed-Form Solution for $\nu=\sqrt{2}$

Abstract
[…]
MSC 2020: 35J05, 78A45, 33E05, 45E10.
Keywords: penetrable wedge diffraction; impedance-matched penetrable wedge; Sommerfeld integrals; Wiener–Hopf; Riemann–Hilbert; elliptic functions; lemniscatic curve; Weierstrass zeta-function.
AI disclosure and responsibility statement
The author used OpenAI’s ChatGPT (GPT-5.2 Pro) as a writing and editing aid for exposition, LaTeX typesetting, and consistency checks. All mathematical derivations, claims, and proofs remain the responsibility of the author.
Intuitive overview
[…]
Roadmap of proof
We write the right-angle penetrable wedge problem as a Sommerfeld spectral representation on a Snell surface that, in the impedance-matched lemniscatic case $\nu=\sqrt{2}$, closes on an elliptic curve. The paper is organized as follows.
  1. 1.
    §1 states the boundary-value problem (Helmholtz transmission, radiation, and Meixner edge condition) and fixes branch and sign conventions.
  2. 2.
    §1.1 introduces the Sommerfeld integral representation and the analytic strip/growth conditions imposed on the spectral densities. We use a standard uniqueness principle for Sommerfeld transforms (Lemma 1.3) to convert equality of boundary integrals into functional relations.
  3. 3.
    §2 derives the lemniscatic Snell surface $\Sigma_{\mathrm{lem}}$ and its Weierstrass uniformization on the square lattice.
  4. 4.
    §4 gives a reproducible prescription for the forcing pole set: each label $\ell=(m,\sigma,\varepsilon_{w})$ determines a spectral point $\zeta_{\ell}$ and hence a point $(t_{\ell},Y_{\ell})\in\Sigma_{\mathrm{lem}}$ and a uniformizing coordinate $u_{\ell}$.
  5. 5.
    §6 records the residue tables $(\alpha_{\ell},\beta_{\ell},C_{\ell})$; their derivation from the global two-face spectral system is given in Appendix A.
  6. 6.
    §7 proves the half-period shift identities needed to eliminate $\wp(u_{0}-u_{\ell})$, $\wp^{\prime}(u_{0}-u_{\ell})$ and $\wp^{\prime\prime}(u_{0}-u_{\ell})$ from the jet coefficients.
  7. 7.
    §8 constructs the jet-killing polynomials $p(t)$ and $q(t)$ and proves the cancellations $A(u(t))+p(t)=O(t^{4})$ and $B(u(t))+q(t)=O(t^{4})$ as $t\to 0$ on the physical component.
  8. 8.
    §10 states the canonical “no double counting” decomposition of $Q_{\mathrm{scat}}$ and proves the pole cancellation properties of the remainder $R(u)$ on the physical cut domain.
  9. 9.
    §11 proves that $Q_{\mathrm{scat}}$ is analytic at the incident spectral point $\zeta=\zeta_{i}$ (limiting absorption), by the explicit exclusion of the incident label from the scattered pole set.
  10. 10.
    §13 derives the far-field diffraction coefficient by steepest descent, under explicit analyticity and nondegeneracy hypotheses.
Notation and conventions
  • $(r,\theta)$ are polar coordinates centered at the wedge apex. The right-angle wedge faces are $\theta=\pm\theta_{w}$ with $\theta_{w}=\pi/4$.
  • The exterior wavenumber is $k_{0}$ and the interior wavenumber is $k_{1}=\nu k_{0}$ with fixed refractive index ratio $\nu=\sqrt{2}$; the impedance match is $\rho=1$.
  • The complex spectral variable is $\zeta\in\mathbb{C}$ (Sommerfeld strip). The incident spectral pole is $\zeta_{i}=\theta_{i}+i\varepsilon$ with $\varepsilon>0$ (limiting absorption), and the physical limit is $\varepsilon\to 0^{+}$.
  • We use the Sommerfeld integration parameter $z$ and the associated variable $t=e^{iz}$.
  • The lemniscatic curve is $\Sigma_{\mathrm{lem}}:=\{(t,Y):\;Y^{2}=2(t^{4}+1)\}$. The physical sheet $\Omega_{+}^{\mathrm{phys}}\subset\Sigma_{\mathrm{lem}}$ is characterized by $|t|<1$ and $Y\to-\sqrt{2}$ as $t\to 0$.
  • $\zeta_{W}(u)$, $\wp(u)$, and $\wp^{\prime}(u)$ denote the Weierstrass zeta and elliptic functions with invariants $(g_{2},g_{3})=(4,0)$ (square lattice $\tau=i$); the subscript distinguishes the Weierstrass zeta function from the spectral variable $\zeta$.
  • The half-period $u_{0}$ is fixed by $\wp(u_{0})=-1$ and $\wp^{\prime}(u_{0})=0$.
  • The Sommerfeld density is decomposed as $Q(\zeta)=Q_{\mathrm{inc}}(\zeta)+Q_{\mathrm{scat}}(\zeta)$ with $Q_{\mathrm{inc}}(\zeta)=(\zeta-\zeta_{i})^{-1}$. Since only differences $Q(\theta+z)-Q(\theta-z)$ enter the field representation, $Q$ is defined up to an additive constant; we fix the gauge by requiring $Q_{\mathrm{scat}}(u_{0})=0$.
1 Introduction and setup
[…]
Scope. The analysis and the resulting closed form apply only to the special configuration $(\theta_{w},\nu,\rho)=(\pi/4,\sqrt{2},1)$. We do not claim an explicit closed-form solution for general penetrable wedges (arbitrary contrast and wedge angle), for which the standard spectral reductions lead to matrix/generalized Wiener–Hopf or multi-variable boundary-value problems; see [4, 7, 9, 10, 12, 11, 13].
[…]
We work in two spatial dimensions and use polar coordinates $(r,\theta)$ about the wedge tip. The penetrable wedge occupies the sector $|\theta|<\theta_{w}$ (medium 1) and is embedded in the exterior $\{|\theta|>\theta_{w}\}$ (medium 0). We consider the scalar Helmholtz transmission problem
$(\Delta+k_{0}^{2})u_{0}=0\quad\text{in }\{|\theta|>\theta_{w}\},\qquad(\Delta+k_{1}^{2})u_{1}=0\quad\text{in }\{|\theta|<\theta_{w}\},$
with $k_{1}=\nu k_{0}$ and refractive index fixed at $\nu=\sqrt{2}$. An incident plane wave in the exterior is
$u_{\mathrm{inc}}(r,\theta)=\exp\!\bigl(ik_{0}r\cos(\theta-\theta_{i})\bigr),$
and we write $u_{0}=u_{\mathrm{inc}}+u_{0,\mathrm{scat}}$ for the total exterior field, while $u_{1}$ denotes the transmitted field. Impedance match ($\rho=1$) reduces the transmission conditions on each face $\theta=\pm\theta_{w}$ to continuity of the field and its normal derivative. Since the unit normal to a radial ray is proportional to $\partial_{\theta}$, these conditions can be written as
$u_{0}(r,\pm\theta_{w})=u_{1}(r,\pm\theta_{w}),\qquad\partial_{\theta}u_{0}(r,\pm\theta_{w})=\partial_{\theta}u_{1}(r,\pm\theta_{w}),\qquad r>0.$
We select the physical solution by the Sommerfeld radiation condition as $r\to\infty$ and the Meixner edge condition at $r=0$ (finite energy near the tip). In the spectral formulation below these requirements are encoded by analyticity and boundedness conditions on the spectral densities.
Let $\theta_{w}=\pi/4$ denote the half-opening angle of the right-angle wedge. We impose limiting absorption by shifting the incident spectral pole off the real axis:
$\zeta_{i}:=\theta_{i}+i\varepsilon,\qquad\varepsilon>0.$ (1)
1.1 Sommerfeld representation and spectral split
A standard Sommerfeld representation of the medium-0 field is
$u^{(0)}(r,\theta)=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{0}r\cos z}\Bigl(Q(\theta+z)-Q(\theta-z)\Bigr)\,dz,$ (2)
for a Sommerfeld contour $\gamma$. The transmitted (medium-1) field admits the analogous representation
$u^{(1)}(r,\theta)=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{1}r\cos z}\Bigl(S(\theta+z)-S(\theta-z)\Bigr)\,dz,\qquad|\theta|<\theta_{w},$ (3)
where $k_{1}=\nu k_{0}$ and $S$ is the medium-1 spectral density. We split
$Q(\zeta)=Q_{\mathrm{inc}}(\zeta)+Q_{\mathrm{scat}}(\zeta),\qquad Q_{\mathrm{inc}}(\zeta)=\frac{1}{\zeta-\zeta_{i}}.$ (4)
Remark 1.1 (Normalization / gauge).
Only the difference $Q(\theta+z)-Q(\theta-z)$ appears in (2). Hence $Q$ is defined up to an additive constant without affecting $u^{(0)}$. We fix this gauge by imposing the normalization
$Q_{\mathrm{scat}}(u_{0})=0,$ (5)
where $u_{0}$ is the half-period point corresponding to $(t,Y)=(0,-\sqrt{2})$ on the physical component (see §3). In the zeta-difference representations used below, the subtraction $\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})$ enforces (5) automatically.
1.2 Scattered allocation
Definition 1.2 (Scattered allocation).
We require
$Q_{\mathrm{scat}}\ \text{is analytic at}\ \zeta=\zeta_{i}.$ (6)
Equivalently, the residue +1+1 at $\zeta=\zeta_{i}$ is carried exclusively by $Q_{\mathrm{inc}}$.
1.3 Transmission conditions in spectral form
The Sommerfeld representations are designed so that the transmission conditions on a face $\theta=\theta_{b}\in\{\pm\theta_{w}\}$ reduce to algebraic relations among boundary values of $Q$ and $S$. Let $w=w_{b}(z)$ denote the Snell map for the face $\theta=\theta_{b}$, defined by matching the oscillatory factors:
$k_{0}\cos w_{b}(z)=k_{1}\cos z=\nu k_{0}\cos z,$ (7)
with the branch determined by $w_{b}(z)\sim z+i\log\nu$ as $\Im z\to+\infty$. Write $w_{b}^{\prime}(z)=dw_{b}/dz$.
Lemma 1.3 (Sommerfeld nullity / uniqueness).
Let $\gamma$ be the Sommerfeld contour and strip described in §1.1. Suppose $H(\zeta)$ is analytic in that strip, satisfies the stated growth/decay bounds along $\gamma$, and define
$U(r,\theta):=\frac{1}{2\pi i}\int_{\gamma}e^{ikr\cos\zeta}\,H(\zeta)\,d\zeta.$
If $U(r,\theta)=0$ for all $r>0$ and for $\theta$ in an interval of length $2\theta_{w}$, then $H(\zeta)\equiv 0$ in the strip.
Proof.
A proof under hypotheses matching the present strip and growth conditions is standard; see, for example, [6, §2] or [4, §2.2]. We invoke this uniqueness principle only in the following form: if two spectral densities produce identical Sommerfeld integrals on a wedge face for all $r>0$, then their difference has vanishing Sommerfeld integral and hence the densities coincide. ∎
Lemma 1.3 is the uniqueness principle underlying Sommerfeld/Malyuzhinets representations: it permits one to infer functional relations between spectral densities from vanishing boundary traces. All spectral identities below that equate integrands from equalities of Sommerfeld integrals are justified by Lemma 1.3.
Proposition 1.4 (Face coupling for $\rho=1$).
Assume $Q$ and $S$ are analytic in a common Sommerfeld strip and have sufficient decay so that integration by parts in $z$ produces no boundary terms. Fix a face $\theta=\theta_{b}$ and let $w=w_{b}(z)$ be as in (7). Then the impedance-matched transmission conditions
$u_{0}(r,\theta_{b})=u_{1}(r,\theta_{b}),\qquad\partial_{\theta}u_{0}(r,\theta_{b})=\partial_{\theta}u_{1}(r,\theta_{b}),\qquad r>0,$
are equivalent to the pointwise spectral relation
$\begin{pmatrix}S(\theta_{b}+z)\\[2.0pt] S(\theta_{b}-z)\end{pmatrix}=\frac{1}{2}\begin{pmatrix}1+w_{b}^{\prime}(z)&1-w_{b}^{\prime}(z)\\[2.0pt] 1-w_{b}^{\prime}(z)&1+w_{b}^{\prime}(z)\end{pmatrix}\begin{pmatrix}Q(\theta_{b}+w_{b}(z))\\[2.0pt] Q(\theta_{b}-w_{b}(z))\end{pmatrix}$. (8)
Proof.
We record the standard argument, included here to make the reduction from the transmission problem explicit. Evaluate (2) and its $\theta$-derivative at $\theta=\theta_{b}$. Differentiating under the integral sign gives
$\partial_{\theta}u_{0}(r,\theta_{b})=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{0}r\cos z}\bigl[Q^{\prime}(\theta_{b}+z)-Q^{\prime}(\theta_{b}-z)\bigr]\,dz.$
Since $Q^{\prime}(\theta_{b}+z)-Q^{\prime}(\theta_{b}-z)=\frac{d}{dz}\bigl(Q(\theta_{b}+z)+Q(\theta_{b}-z)\bigr)$, an integration by parts yields
$\partial_{\theta}u_{0}(r,\theta_{b})=\frac{k_{0}r}{2\pi}\int_{\gamma}\sin z\,e^{ik_{0}r\cos z}\bigl[Q(\theta_{b}+z)+Q(\theta_{b}-z)\bigr]\,dz.$
An identical computation for $u_{1}$ gives
$\partial_{\theta}u_{1}(r,\theta_{b})=\frac{k_{1}r}{2\pi}\int_{\gamma}\sin z\,e^{ik_{1}r\cos z}\bigl[S(\theta_{b}+z)+S(\theta_{b}-z)\bigr]\,dz.$
Now change variables in the medium-1 integrals by $w=w_{b}(z)$, so that $e^{ik_{1}r\cos z}=e^{ik_{0}r\cos w}$ by (7). The change of variables gives $dz=dw/w_{b}^{\prime}(z)$, while differentiating (7) implies $\sin z=\sin w\,w_{b}^{\prime}(z)/\nu$. Using $k_{1}=\nu k_{0}$, we obtain
$u_{1}(r,\theta_{b})$ $=\frac{1}{2\pi i}\int e^{ik_{0}r\cos w}\,\frac{S(\theta_{b}+z)-S(\theta_{b}-z)}{w_{b}^{\prime}(z)}\,dw,$
$\partial_{\theta}u_{1}(r,\theta_{b})$ $=\frac{k_{0}r}{2\pi}\int\sin w\,e^{ik_{0}r\cos w}\bigl[S(\theta_{b}+z)+S(\theta_{b}-z)\bigr]\,dw,$
with $z=z(w)$ the inverse map. Comparing with the corresponding expressions for $u_{0}$ and $\partial_{\theta}u_{0}$, and using Lemma 1.3, we obtain the two scalar identities
$\frac{S(\theta_{b}+z)-S(\theta_{b}-z)}{w_{b}^{\prime}(z)}=Q(\theta_{b}+w)-Q(\theta_{b}-w),\qquad S(\theta_{b}+z)+S(\theta_{b}-z)=Q(\theta_{b}+w)+Q(\theta_{b}-w),$
which solve to (8). ∎
Theorem 1.5 (Main results for the lemniscatic right-angle wedge).
Assume the impedance-matched right-angle configuration $\theta_{w}=\pi/4$, $\nu=\sqrt{2}$, $\rho=1$, and $\varepsilon>0$. Let
$Q_{\mathrm{inc}}(\zeta)=\frac{1}{\zeta-(\theta_{i}+i\varepsilon)},\qquad Q(\zeta)=Q_{\mathrm{inc}}(\zeta)+Q_{\mathrm{scat}}(\zeta),$
and define the physical branch lift $\zeta\mapsto u(\zeta)$ by the lemniscatic Snell surface (Section 2) and the Weierstrass uniformization (Section 3). Then:
  1. (i)
    (Canonical representation.) The scattered spectral density $Q_{\mathrm{scat}}$ admits the decomposition
    $Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\bigr]+R(u),$
    where the pole set $I_{\mathrm{scat}}$ and points $u_{\ell}$ are defined in Section 4, the coefficients $C_{\ell}$ are given explicitly in Proposition 6.3, and the remainder $R$ is pole-free at each forcing pole $u_{\ell}$ and analytic at $u_{0}$ (Theorem 10.4).
  2. (ii)
    (Incident analyticity.) The scattered part is analytic at the incident spectral point $\zeta=\theta_{i}+i\varepsilon$ (Theorem 11.1).
  3. (iii)
    (Far-field coefficient.) The diffracted far-field coefficient in medium 0 is
    $D(\theta,\theta_{i})=e^{-i3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$
    where $Q_{\mathrm{scat}}(\theta)$ denotes the physical branch boundary value of $Q_{\mathrm{scat}}(\zeta)$ at $\zeta=\theta$ (Theorem 13.1).
2 Lemniscatic Snell surface and physical branch
2.1 Lemniscatic curve
The lemniscatic Snell surface is the algebraic curve
$\Sigma_{\mathrm{lem}}:\qquad Y^{2}=2(t^{4}+1).$ (9)
We work on the physical plus component $\Omega_{+}^{\mathrm{phys}}\subset\Sigma_{\mathrm{lem}}$ characterized by
$|t|<1,\qquad Y\to-\sqrt{2}\ \text{as}\ t\to 0.$ (10)
Origin of the Snell surface in the impedance-matched case. The Sommerfeld representations in the two media use the plane-wave factors $e^{ik_{0}r\cos z}$ (exterior) and $e^{ik_{1}r\cos z}$ (interior). On a wedge face $\theta=\theta_{b}$ one matches these factors by the analytic change of variables $z\mapsto w=w_{b}(z)$ determined implicitly by
$k_{0}\cos w_{b}(z)=k_{1}\cos z=\nu k_{0}\cos z.$
Thus the Snell map $w=w_{b}(z)$ satisfies $\cos w=\nu\cos z$ and is fixed by the radiation/limiting-absorption branch condition $w(z)\sim z+i\log\nu$ as $\Im z\to+\infty$. Writing $t=e^{iz}$ and $s=e^{iw}$, the identity $\cos w=\nu\cos z$ becomes
$\frac{1}{2}\Bigl(s+\frac{1}{s}\Bigr)=\nu\,\frac{1}{2}\Bigl(t+\frac{1}{t}\Bigr),\qquad\text{i.e.}\qquad s^{2}-\nu\Bigl(t+\frac{1}{t}\Bigr)s+1=0.$
In the special lemniscatic case $\nu=\sqrt{2}$ we define
$Y:=2ts-\sqrt{2}\,(t^{2}+1),$
and a short calculation shows that the quadratic relation above is equivalent to $Y^{2}=2(t^{4}+1)$. The physical component $\Omega_{+}^{\mathrm{phys}}$ corresponds to the branch $|t|<1$ with $Y\to-\sqrt{2}$ as $t\to 0$, for which $s\sim t/\sqrt{2}$.
Quarter-period symmetry and orbit branches. The lemniscatic curve admits the order-four automorphism
$\tau:\Sigma_{\mathrm{lem}}\to\Sigma_{\mathrm{lem}},\qquad\tau(t,Y)=(it,-Y),$ (11)
which preserves $\Omega_{+}^{\mathrm{phys}}$. Let $w=w(t,Y)$ denote the (multi-valued) analytic function on $\Sigma_{\mathrm{lem}}$ defined by $e^{iw}=s(t,Y)$, with the physical branch fixed by $w(z)\sim z+i\log\nu$ as $\Im z\to+\infty$ (equivalently $s\sim t/\nu$ as $t\to 0$). Following the standard orbit construction for a right-angle wedge, we introduce four orbit branches $w_{m}$ by
$e^{iw_{m}(t,Y)}:=\bigl(s(\tau^{m}(t,Y))\bigr)^{(-1)^{m}},\qquad m\in\{0,1,2,3\}.$ (12)
The forcing poles are transported along these orbits and labelled by $(m,\sigma,\varepsilon_{w})$ in Section 4.
2.2 Physical Snell exponential and spectral map
Define the physical Snell exponential on $\Sigma_{\mathrm{lem}}$ by
$s(t,Y):=\frac{\sqrt{2}(t^{2}+1)+Y}{2t}.$ (13)
Differentiating $\cos w=\nu\cos z$ with $e^{iw}=s(t,Y)$ and $t=e^{iz}$ yields the algebraic derivative
$g^{\prime}(t,Y)=\frac{dw}{dz}=\frac{\sqrt{2}(t^{2}-1)}{Y}.$ (14)
and the spectral exponential
$s_{\zeta}:=\exp\bigl(i(\zeta-\theta_{w})\bigr).$ (15)
For $\zeta$ in the Sommerfeld strip (with $\varepsilon>0$ fixed), the physical branch map $\zeta\mapsto(t(\zeta),Y(\zeta))\in\Omega_{+}^{\mathrm{phys}}$ is defined by solving
$s(t(\zeta),Y(\zeta))=s_{\zeta},\qquad|t(\zeta)|<1,\qquad Y(\zeta)\to-\sqrt{2}\ \text{as}\ t(\zeta)\to 0.$ (16)
3 Weierstrass uniformization for $\tau=i$
We take the square lattice $\tau=i$ with Weierstrass invariants
$(g_{2},g_{3})=(4,0).$ (17)
Let $\wp(u)$, $\wp^{\prime}(u)$ and $\zeta_{W}(u)$ denote the corresponding Weierstrass elliptic and zeta functions; see, e.g., [14, 15, 16]. We adopt the uniformization
$\wp(u)=x_{W}(t):=\frac{it^{2}+1}{it^{2}-1},\qquad\frac{1}{2}\,\wp^{\prime}(u)=y_{W}(t,Y):=\frac{(i-1)tY}{(it^{2}-1)^{2}}.$ (18)
The physical lift $u=u(\zeta)$ is selected by composing the physical branch $\zeta\mapsto(t(\zeta),Y(\zeta))$ with the uniformization on $\Omega_{+}^{\mathrm{phys}}$.
Let $u_{0}$ denote the half-period corresponding to the physical point $(t,Y)=(0,-\sqrt{2})$, so that
$\wp(u_{0})=-1,\qquad\wp^{\prime}(u_{0})=0.$ (19)
4 Pole set, labels, and incident exclusion
Poles are indexed by
$\ell=(m,\sigma,\varepsilon_{w}),\qquad m\in\{0,1,2,3\},\quad\sigma\in\{\pm 1\},\quad\varepsilon_{w}\in\{\pm 1\}.$ (20)
Define the map $(\sigma,\varepsilon_{w})\mapsto j$ by
$(+,+)\mapsto 3,\qquad(+,-)\mapsto 1,\qquad(-,+)\mapsto 4,\qquad(-,-)\mapsto 2,$ (21)
and the sign
$\varepsilon_{j}=\begin{cases}+1,&j\in\{1,3\},\\ -1,&j\in\{2,4\}.\end{cases}$ (22)
The incident label is $\ell_{\mathrm{inc}}=(0,+,-)$ and the scattered index set is
$I_{\mathrm{scat}}:=I\setminus\{\ell_{\mathrm{inc}}\},\qquad|I_{\mathrm{scat}}|=15,$ (23)
where $I=\{0,1,2,3\}\times\{\pm 1\}\times\{\pm 1\}$.
For later reference we make the pole label $\ell\mapsto(t_{\ell},Y_{\ell})\mapsto u_{\ell}$ explicit. Set the limiting-absorption incident angle $\zeta_{i}=\theta_{i}+i\varepsilon$ and fix $\theta_{w}=\pi/4$. For $(\sigma,\varepsilon_{w})\in\{\pm 1\}^{2}$ define the four forcing phases
$a_{\sigma,\varepsilon_{w}}:=\exp\bigl(i\sigma(\zeta_{i}+\varepsilon_{w}\theta_{w})\bigr),\qquad b_{m,\sigma,\varepsilon_{w}}:=a_{\sigma,\varepsilon_{w}}^{(-1)^{m}}\quad(m=0,1,2,3).$ (24)
The physical orbit table fixes the pole condition in the form
$e^{iw_{m}}=(s(\tau^{m}p))^{(-1)^{m}}=b_{m,\sigma,\varepsilon_{w}},$
so that the forcing pole points are the solutions of $s(q)=b$ on $\Sigma_{\mathrm{lem}}$ with $q=\tau^{m}p$.
Lemma 4.1 (Explicit algebraic pole points on $\Sigma_{\mathrm{lem}}$).
Fix $b\in\mathbb{C}\setminus\{0\}$ and consider the equation $s(t,Y)=b$ on $\Sigma_{\mathrm{lem}}$, with $s$ defined by (13). Then $Y$ is forced to be
$Y=2bt-\sqrt{2}\,(t^{2}+1),$ (25)
and $(t,Y)\in\Sigma_{\mathrm{lem}}$ if and only if $t$ satisfies the quadratic
$t^{2}-\frac{b^{2}+1}{\sqrt{2}\,b}\,t+1=0.$ (26)
Equivalently, the two roots are
$t_{\pm}(b)=\frac{(b^{2}+1)\pm\sqrt{b^{4}-6b^{2}+1}}{2\sqrt{2}\,b},\qquad t_{+}(b)\,t_{-}(b)=1.$ (27)
On the physical component $\Omega_{+}^{\mathrm{phys}}$ one selects the root $t_{\mathrm{in}}(b)\in\mathbb{D}:=\{|t|<1\}$, and the other root is $t_{\mathrm{out}}(b)=1/t_{\mathrm{in}}(b)$.
Proof.
Starting from $s(t,Y)=b$ and (13), we solve for $Y$ to obtain (25). Substituting (25) into $Y^{2}=2(t^{4}+1)$ yields
$\bigl(2bt-\sqrt{2}(t^{2}+1)\bigr)^{2}=2(t^{4}+1),$
which simplifies to $b^{2}+1-\sqrt{2}\,b(t+1/t)=0$ after division by $2t^{2}$. Multiplying by $t$ gives (26), and the quadratic formula yields (27). The product identity $t_{+}t_{-}=1$ is immediate from (26). The physical selection $|t|<1$ defines $t_{\mathrm{in}}$. ∎
For $\ell=(m,\sigma,\varepsilon_{w})$ we set $b=b_{m,\sigma,\varepsilon_{w}}$ and define the intermediate point $q$ by
$t_{q}:=t_{\mathrm{in}}(b),\qquad Y_{q}:=2b\,t_{q}-\sqrt{2}\,(t_{q}^{2}+1).$
Transporting back to the base point $p=\tau^{-m}q$ gives the pole coordinates on $\Sigma_{\mathrm{lem}}$:
$(t_{\ell},Y_{\ell})=\bigl(i^{-m}t_{q},\,(-1)^{m}Y_{q}\bigr).$ (28)
Finally, the corresponding lift $u_{\ell}$ on the uniformizing torus is defined by the Weierstrass map
$\wp(u_{\ell})=x_{W}(t_{\ell}),\qquad\frac{1}{2}\,\wp^{\prime}(u_{\ell})=y_{W}(t_{\ell},Y_{\ell}),$ (29)
with the physical lift selected on $\Omega_{+}^{\mathrm{phys}}$ (equivalently, by fixing the sign of $\wp^{\prime}(u_{\ell})$ through $y_{W}(t_{\ell},Y_{\ell})$).
5 Derivative/residue conventions and phase symbols
5.1 Derivative and residue conventions
We adopt the orbit derivative
$w_{m}^{\prime}(u)=\frac{\,\mathrm{d}w_{m}}{\,\mathrm{d}u}=\begin{cases}-e^{-i\pi/4}\Bigl(t-\dfrac{1}{t}\Bigr),&m\ \text{even},\\[4.30554pt] +e^{-i\pi/4}\Bigl(t+\dfrac{1}{t}\Bigr),&m\ \text{odd},\end{cases}\qquad t=t(u).$ (30)
Lemma 5.1.
The orbit derivatives in (30) follow from the lemniscatic Snell relation (13) and the uniformization (18).
Proof.
Write $w=w_{0}$ for the physical branch defined by $e^{iw}=s(t,Y)$, and set $t=t(u)$, $Y=Y(u)$ along the physical lift. Differentiating (18) gives
$\wp^{\prime}(u)=\frac{d}{du}x_{W}\bigl(t(u)\bigr)=x_{W}^{\prime}(t)\,t^{\prime}(u).$
A direct computation from $x_{W}(t)=(it^{2}+1)/(it^{2}-1)$ yields
$x_{W}^{\prime}(t)=-\frac{4it}{(it^{2}-1)^{2}},$
while $\wp^{\prime}(u)=2y_{W}(t,Y)=2(i-1)tY/(it^{2}-1)^{2}$. Equating these expressions shows
$\frac{dt}{du}=t^{\prime}(u)=-\frac{1+i}{2}\,Y(u).$ (31)
Next, differentiating the Snell identity $\cos w=\nu\cos z$ with $\nu=\sqrt{2}$ and using $t=e^{iz}$ gives
$\sin w\,\frac{dw}{dz}=-\sqrt{2}\,\sin z\qquad\Rightarrow\qquad\frac{dw}{dz}=\sqrt{2}\,\frac{\sin z}{\sin w}.$
Expressing $\sin z=(t-1/t)/(2i)$ and $\sin w=(s-1/s)/(2i)$, and using the relation $s=s(t,Y)$, one obtains the algebraic form already recorded in (14):
$\frac{dw}{dz}=g^{\prime}(t,Y)=\sqrt{2}\,\frac{t^{2}-1}{Y}.$
Since $dt/dz=it$, we have
$\frac{dw}{dt}=\frac{dw/dz}{dt/dz}=-i\sqrt{2}\,\frac{t^{2}-1}{tY}.$
Combining this with (31) gives
$\frac{dw}{du}=\frac{dw}{dt}\frac{dt}{du}=-i\sqrt{2}\,\frac{t^{2}-1}{tY}\Bigl(-\frac{1+i}{2}Y\Bigr)=-e^{-i\pi/4}\Bigl(t-\frac{1}{t}\Bigr),$
which is the $m$ even case in (30). For the orbit branches $w_{m}$ defined by (12), the additional factors $\tau^{m}$ and $(-1)^{m}$ change $t$ to $i^{m}t$ and introduce the sign change for odd $m$, yielding the $m$ odd case in (30). ∎
Define
$r_{I}(\ell):=\frac{\varepsilon_{j}}{w_{m}^{\prime}(u_{\ell})}.$ (32)
Equivalently, the explicit forms are
$r_{I}(\ell)=\begin{cases}-\varepsilon_{j}\,e^{i\pi/4}\dfrac{t_{\ell}}{t_{\ell}^{2}-1},&m\ \text{even},\\[5.16663pt] +\varepsilon_{j}\,e^{i\pi/4}\dfrac{t_{\ell}}{t_{\ell}^{2}+1},&m\ \text{odd}.\end{cases}$ (33)
5.2 Phase symbols
We use the phase symbols
$\chi_{m}:=i^{m}\in\{\pm 1\}\quad(m\ \text{even}),\qquad\psi_{m}:=i^{m}\in\{\pm i\},\qquad\kappa_{m}:=i^{m+1}\in\{\pm 1\}\quad(m\ \text{odd}).$ (34)
6 Residue data and per-pole jet summands
The coefficients $(\alpha_{\ell})$, $(\beta_{\ell})$ and $(C_{\ell})$ appearing in the elliptic reconstruction are obtained by solving the global two-face spectral functional system and evaluating the forcing residues at each pole $u_{\ell}$. For readability we record the resulting closed-form tables below; the derivation is given in Appendix A.
6.1 Residue data: $(\alpha_{\ell})$, $(\beta_{\ell})$, $(C_{\ell})$
Throughout this section, for a fixed pole label $\ell$ we write $(t,Y)=(t_{\ell},Y_{\ell})$ and $r_{I}=r_{I}(\ell)$.
Proposition 6.1 (Coefficient table for $(\alpha_{\ell})$).
For $\ell\in I_{\mathrm{scat}}$ with $j=j(\sigma,\varepsilon_{w})$:
$\text{if $m$ is even:}\quad\alpha_{\ell}$ $=\begin{cases}-\chi_{m}\,r_{I}\,t^{2},&j=1,\\ -\chi_{m}\,r_{I}\,(t^{4}-t^{2}+1),&j=2,\\ 0,&j=3,\\ -\chi_{m}\,r_{I}\,\dfrac{iY}{\sqrt{2}}(t^{2}-1),&j=4;\end{cases}$ (35)
$\text{if $m$ is odd:}\quad\alpha_{\ell}$ $=\begin{cases}\psi_{m}\,r_{I}\,t^{4},&j=1,\\ \psi_{m}\,r_{I},&j=2,\\ \kappa_{m}\,r_{I}\,\dfrac{Y}{\sqrt{2}}\,t^{2},&j=3,\\ -\kappa_{m}\,r_{I}\,\dfrac{Y}{\sqrt{2}},&j=4.\end{cases}$ (36)
Proof.
See Appendix A. ∎
Proposition 6.2 (Coefficient table for $(\beta_{\ell})$).
For $\ell\in I_{\mathrm{scat}}$ with $j=j(\sigma,\varepsilon_{w})$:
$\text{if $m$ is even:}\quad\beta_{\ell}$ $=\begin{cases}\chi_{m}\,r_{I}\,\dfrac{i}{\sqrt{2}}Y\,t^{2},&j=1,\\ \chi_{m}\,r_{I}\,\dfrac{i}{\sqrt{2}}Y,&j=2,\\ -\chi_{m}\,r_{I}\,t^{4},&j=3,\\ -\chi_{m}\,r_{I},&j=4;\end{cases}$ (37)
$\text{if $m$ is odd:}\quad\beta_{\ell}$ $=\begin{cases}0,&j=1,\\ \kappa_{m}\,r_{I}\,\dfrac{1}{\sqrt{2}}Y\,(t^{2}+1),&j=2,\\ \psi_{m}\,r_{I}\,t^{2},&j=3,\\ -\psi_{m}\,r_{I}\,(t^{4}+t^{2}+1),&j=4.\end{cases}$ (38)
Proof.
See Appendix A. ∎
Proposition 6.3 (Global residues $(C_{\ell})$).
For $\ell\in I_{\mathrm{scat}}$ with $j=j(\sigma,\varepsilon_{w})$:
$\text{if $m$ is even:}\quad C_{\ell}$ $=\begin{cases}-\chi_{m}\,\dfrac{r_{I}}{2t^{2}},&j=1,\\[3.87495pt] -\dfrac{r_{I}}{2}\,\dfrac{2t^{4}-2t^{2}+1}{t^{4}},&j=2,\ m=0,\\[5.16663pt] +\dfrac{r_{I}}{2}\,\dfrac{1}{t^{4}},&j=2,\ m=2,\\[3.87495pt] 0,&j=3,\\ 0,&j=4;\end{cases}$ (39)
$\text{if $m$ is odd:}\quad C_{\ell}$ $=\begin{cases}0,&j=1,\\ 0,&j=2,\\[3.44444pt] \dfrac{r_{I}}{2}\Bigl(1+\kappa_{m}\,\dfrac{Y}{\sqrt{2}\,t^{2}}\Bigr),&j=3,\\[5.16663pt] -\dfrac{r_{I}}{2t^{2}}\Bigl(1+\kappa_{m}\,\dfrac{Y}{\sqrt{2}\,t^{2}}\Bigr),&j=4.\end{cases}$ (40)
Proof.
See Appendix A. ∎
6.2 Per-pole jet summands
We define jet polynomials
$p(t)=p_{1}t+p_{2}t^{2}+p_{3}t^{3},\qquad q(t)=q_{1}t+q_{2}t^{2}+q_{3}t^{3},$ (41)
with coefficients $p_{n}=\sum_{\ell\in I_{\mathrm{scat}}}p_{n}^{(\ell)}$ and $q_{n}=\sum_{\ell\in I_{\mathrm{scat}}}q_{n}^{(\ell)}$. For later reference we record the per-pole contributions to the jet-killing coefficients in closed form.
In this section we record the explicit coefficient formulas; their derivation is given in Appendix A.
Proposition 6.4 (Per-pole $p$-summands).
Evaluate at $(t,Y)=(t_{\ell},Y_{\ell})$.
  • $m$ even ($\chi_{m}$).
    $j=1:\$$p_{1}^{(\ell)}=\chi_{m}\,i\,\frac{t}{t^{2}-1},\qquad p_{2}^{(\ell)}=-\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Y}{t^{2}-1},\qquad p_{3}^{(\ell)}=\chi_{m}\,i\,\frac{t^{4}+3}{3t(t^{2}-1)}.$
    $j=2:\$$p_{1}^{(\ell)}=\chi_{m}\,i\,\frac{-t^{4}+t^{2}-1}{t(t^{2}-1)},\qquad p_{2}^{(\ell)}=+\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Y(t^{4}-t^{2}+1)}{t^{2}(t^{2}-1)},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=-\chi_{m}\,i\,\frac{(t^{4}+3)(t^{4}-t^{2}+1)}{3t^{3}(t^{2}-1)}.$
    $j=3:\$$p_{1}^{(\ell)}=0,\qquad p_{2}^{(\ell)}=0,\qquad p_{3}^{(\ell)}=0.$
    $j=4:\$$p_{1}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}\,t},\qquad p_{2}^{(\ell)}=-\chi_{m}\,\frac{t^{4}+1}{t^{2}},\qquad p_{3}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t^{4}+3}{3t^{3}}.$
  • $m$ odd ($\psi_{m},\kappa_{m}$).
    $j=1:\$$p_{1}^{(\ell)}=i\,\psi_{m}\,\frac{t^{3}}{t^{2}+1},\qquad p_{2}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Yt^{2}}{t^{2}+1},\qquad p_{3}^{(\ell)}=i\,\psi_{m}\,\frac{t(t^{4}+3)}{3(t^{2}+1)}.$
    $j=2:\$$p_{1}^{(\ell)}=-i\,\psi_{m}\,\frac{1}{t(t^{2}+1)},\qquad p_{2}^{(\ell)}=+\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Y}{t^{2}(t^{2}+1)},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=-i\,\psi_{m}\,\frac{t^{4}+3}{3t^{3}(t^{2}+1)}.$
    $j=3:\$$p_{1}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Yt}{t^{2}+1},\qquad p_{2}^{(\ell)}=-i\,\kappa_{m}\,\frac{t^{4}+1}{t^{2}+1},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y(t^{4}+3)}{3t(t^{2}+1)}.$
    $j=4:\$$p_{1}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y}{t(t^{2}+1)},\qquad p_{2}^{(\ell)}=-i\,\kappa_{m}\,\frac{t^{4}+1}{t^{2}(t^{2}+1)},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y(t^{4}+3)}{3t^{3}(t^{2}+1)}.$
Proposition 6.5 (Per-pole $q$-summands).
Evaluate at $(t,Y)=(t_{\ell},Y_{\ell})$.
  • $m$ even ($\chi_{m}$).
    $j=1:\$$q_{1}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t}{t^{2}-1},\qquad q_{2}^{(\ell)}=-\chi_{m}\,\frac{t^{4}+1}{t^{2}-1},\qquad q_{3}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t^{4}+3}{3t(t^{2}-1)}.$
    $j=2:\$$q_{1}^{(\ell)}=-\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{1}{t(t^{2}-1)},\qquad q_{2}^{(\ell)}=+\chi_{m}\,\frac{t^{4}+1}{t^{2}(t^{2}-1)},$
    $\hskip 45.00006ptq_{3}^{(\ell)}=-\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t^{4}+3}{3t^{3}(t^{2}-1)}.$
    $j=3:\$$q_{1}^{(\ell)}=\chi_{m}\,i\,\frac{t^{3}}{t^{2}-1},\qquad q_{2}^{(\ell)}=-\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Yt^{2}}{t^{2}-1},\qquad q_{3}^{(\ell)}=\chi_{m}\,i\,\frac{t(t^{4}+3)}{3(t^{2}-1)}.$
    $j=4:\$$q_{1}^{(\ell)}=-\chi_{m}\,i\,\frac{1}{t(t^{2}-1)},\qquad q_{2}^{(\ell)}=+\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Y}{t^{2}(t^{2}-1)},\qquad q_{3}^{(\ell)}=-\chi_{m}\,i\,\frac{t^{4}+3}{3t^{3}(t^{2}-1)}.$
  • $m$ odd ($\psi_{m},\kappa_{m}$).
    $j=1:\$$q_{1}^{(\ell)}=0,\qquad q_{2}^{(\ell)}=0,\qquad q_{3}^{(\ell)}=0.$
    $j=2:\$$q_{1}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y}{t},\qquad q_{2}^{(\ell)}=i\,\kappa_{m}\,\frac{t^{4}+1}{t^{2}},\qquad q_{3}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y(t^{4}+3)}{3t^{3}}.$
    $j=3:\$$q_{1}^{(\ell)}=i\,\psi_{m}\,\frac{t}{t^{2}+1},\qquad q_{2}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Y}{t^{2}+1},\qquad q_{3}^{(\ell)}=i\,\psi_{m}\,\frac{t^{4}+3}{3t(t^{2}+1)}.$
    $j=4:\$$q_{1}^{(\ell)}=i\,\psi_{m}\,\frac{t^{4}+t^{2}+1}{t(t^{2}+1)},\qquad q_{2}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Y(t^{4}+t^{2}+1)}{t^{2}(t^{2}+1)},$
    $\hskip 45.00006ptq_{3}^{(\ell)}=i\,\psi_{m}\,\frac{(t^{4}+3)(t^{4}+t^{2}+1)}{3t^{3}(t^{2}+1)}.$
Remark 6.6.
In the case $m$ odd and $j=3$, the coefficient $q_{3}^{(\ell)}$ contains no factor of $Y$.
7 Half-period shift identities at $e_{2}=-1$
We prove the identities
$W0_{\ell}:=\wp(u_{0}-u_{\ell})=\frac{i}{t_{\ell}^{2}},\qquad W1_{\ell}:=\wp^{\prime}(u_{0}-u_{\ell})=\frac{(1-i)Y_{\ell}}{t_{\ell}^{3}},\qquad W2_{\ell}:=\wp^{\prime\prime}(u_{0}-u_{\ell})=-2-\frac{6}{t_{\ell}^{4}}.$ (42)
7.1 Specialization of the addition theorem
Start from the general addition theorem (see, e.g., [14, § 23.10] or [15, Ch. 20])
$\wp(u+v)=-\wp(u)-\wp(v)+\frac{1}{4}\left(\frac{\wp^{\prime}(u)-\wp^{\prime}(v)}{\wp(u)-\wp(v)}\right)^{2}.$ (43)
Specialize to $v=u_{0}$ with $\wp(u_{0})=-1$ and $\wp^{\prime}(u_{0})=0$:
$\wp(u+u_{0})=-\wp(u)+1+\frac{1}{4}\left(\frac{\wp^{\prime}(u)}{\wp(u)+1}\right)^{2}.$ (44)
Using the differential equation for $(g_{2},g_{3})=(4,0)$ (see, e.g., [14, § 23.6]),
$(\wp^{\prime}(u))^{2}=4\wp(u)^{3}-4\wp(u)=4\wp(u)\bigl(\wp(u)^{2}-1\bigr),$ (45)
we obtain
$\frac{1}{4}\left(\frac{\wp^{\prime}(u)}{\wp(u)+1}\right)^{2}=\frac{\wp(u)\left(\wp(u)^{2}-1\right)}{(\wp(u)+1)^{2}}=\frac{\wp(u)(\wp(u)-1)}{\wp(u)+1}.$ (46)
Substituting yields the half-period shift identity
$\boxed{\ \wp(u+u_{0})=-1+\frac{2}{\wp(u)+1}\ }.$ (47)
Differentiating (47) gives
$\boxed{\ \wp^{\prime}(u+u_{0})=-\frac{2\wp^{\prime}(u)}{(\wp(u)+1)^{2}}\ }.$ (48)
Differentiating (48) yields
$\boxed{\ \wp^{\prime\prime}(u+u_{0})=-2\left(\frac{\wp^{\prime\prime}(u)}{(\wp(u)+1)^{2}}-\frac{2\wp^{\prime}(u)^{2}}{(\wp(u)+1)^{3}}\right)\ }.$ (49)
7.2 Specialization to $u_{0}-u_{\ell}$ and algebraic elimination via the uniformization
Set $u=-u_{\ell}$. Using even/oddness of $\wp$ and $\wp^{\prime}$,
$\wp(-u_{\ell})=\wp(u_{\ell}),\qquad\wp^{\prime}(-u_{\ell})=-\wp^{\prime}(u_{\ell}),\qquad\wp^{\prime\prime}(-u_{\ell})=\wp^{\prime\prime}(u_{\ell}),$ (50)
we obtain from (47)–(48)
$\wp(u_{0}-u_{\ell})=-1+\frac{2}{\wp(u_{\ell})+1}=\frac{1-\wp(u_{\ell})}{1+\wp(u_{\ell})},\qquad\wp^{\prime}(u_{0}-u_{\ell})=\frac{2\wp^{\prime}(u_{\ell})}{(\wp(u_{\ell})+1)^{2}}.$ (51)
Now substitute the uniformization $\wp(u_{\ell})=x_{W}(t_{\ell})$ and $\wp^{\prime}(u_{\ell})=2y_{W}(t_{\ell},Y_{\ell})$. A direct simplification gives
$\wp(u_{0}-u_{\ell})=\frac{i}{t_{\ell}^{2}},\qquad\wp^{\prime}(u_{0}-u_{\ell})=\frac{(1-i)Y_{\ell}}{t_{\ell}^{3}}.$ (52)
Finally, using $\wp^{\prime\prime}(u)=6\wp(u)^{2}-g_{2}/2=6\wp(u)^{2}-2$ for $g_{2}=4$,
$\wp^{\prime\prime}(u_{0}-u_{\ell})=6\left(\frac{i}{t_{\ell}^{2}}\right)^{2}-2=-2-\frac{6}{t_{\ell}^{4}},$ (53)
which proves (42).
8 Jet-killing construction and jet cancellation
8.1 Definitions
Let $\zeta_{W}(u)$ denote the Weierstrass zeta function, characterized by $\zeta_{W}^{\prime}(u)=-\wp(u)$ and $\zeta_{W}(u)\sim u^{-1}$ as $u\to 0$; see [14, § 23.2,§ 23.9] or [16, Ch. 20]. Define
$A(u)=\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr],\qquad B(u)=\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr].$ (54)
Define jet-killing polynomials
$p(t)=p_{1}t+p_{2}t^{2}+p_{3}t^{3},\qquad q(t)=q_{1}t+q_{2}t^{2}+q_{3}t^{3},$ (55)
with coefficients fixed by
$p_{1}$ $=e^{-i\pi/4}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\,W0_{\ell},\qquad p_{2}=e^{-i\pi/2}\frac{1}{2}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\,W1_{\ell},\qquad p_{3}=e^{-3i\pi/4}\frac{1}{6}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\,W2_{\ell},$ (56)
$q_{1}$ $=e^{-i\pi/4}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\,W0_{\ell},\qquad q_{2}=e^{-i\pi/2}\frac{1}{2}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\,W1_{\ell},\qquad q_{3}=e^{-3i\pi/4}\frac{1}{6}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\,W2_{\ell}.$ (57)
8.2 Local relation $u-u_{0}=e^{-i\pi/4}t+O(t^{5})$ (coefficient matching)
Let $\delta:=u-u_{0}$. From the uniformization $\wp(u)=x_{W}(t)$, expand $x_{W}(t)$ as $t\to 0$:
$x_{W}(t)=\frac{it^{2}+1}{it^{2}-1}=-1-2it^{2}+2t^{4}+O(t^{6}).$ (58)
Next expand $\wp(u_{0}+\delta)$. Since $\wp^{\prime}(u_{0})=0$, only even powers appear:
$\wp(u_{0}+\delta)=\wp(u_{0})+\frac{\wp^{\prime\prime}(u_{0})}{2}\delta^{2}+\frac{\wp^{(4)}(u_{0})}{24}\delta^{4}+O(\delta^{6}).$ (59)
With $\wp(u_{0})=-1$ and $\wp^{\prime\prime}(u)=6\wp(u)^{2}-2$, we have $\wp^{\prime\prime}(u_{0})=4$. Moreover, $\wp^{(3)}(u)=12\wp(u)\wp^{\prime}(u)$ so $\wp^{(3)}(u_{0})=0$, and $\wp^{(4)}(u)=12(\wp^{\prime}(u))^{2}+12\wp(u)\wp^{\prime\prime}(u)$ gives $\wp^{(4)}(u_{0})=12(-1)\cdot 4=-48$. Thus
$\wp(u_{0}+\delta)=-1+2\delta^{2}-2\delta^{4}+O(\delta^{6}).$ (60)
Equating with $x_{W}(t)$ yields
$2\delta^{2}-2\delta^{4}=-2it^{2}+2t^{4}+O(t^{6}).$ (61)
Write $\delta^{2}=-it^{2}+at^{4}+O(t^{6})$ so that $\delta^{4}=(-it^{2})^{2}+O(t^{6})=-t^{4}+O(t^{6})$. Matching coefficients forces $a=0$, hence
$\delta=e^{-i\pi/4}t+O(t^{5}),$ (62)
which explicitly rules out any $t^{3}$ term.
8.3 Jet cancellation
Fix $\ell$. Taylor expand $\zeta_{W}$ about $u_{0}-u_{\ell}$ using $\zeta_{W}^{\prime}=$-$\wp$:
$\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})=-\delta\,\wp(u_{0}-u_{\ell})-\frac{\delta^{2}}{2}\wp^{\prime}(u_{0}-u_{\ell})-\frac{\delta^{3}}{6}\wp^{\prime\prime}(u_{0}-u_{\ell})+O(\delta^{4}),$ (63)
where $\delta=u-u_{0}$. Multiply by $\alpha_{\ell}$ and sum over $\ell\in I_{\mathrm{scat}}$ to obtain
$A(u)=-\delta\sum_{\ell}\alpha_{\ell}W0_{\ell}-\frac{\delta^{2}}{2}\sum_{\ell}\alpha_{\ell}W1_{\ell}-\frac{\delta^{3}}{6}\sum_{\ell}\alpha_{\ell}W2_{\ell}+O(\delta^{4}).$ (64)
Using (62) gives $\delta^{n}=e^{-in\pi/4}t^{n}+O(t^{n+4})$ for $n=1,2,3$, hence
$A(u(t))$ $=-e^{-i\pi/4}t\sum_{\ell}\alpha_{\ell}W0_{\ell}-e^{-i\pi/2}\frac{t^{2}}{2}\sum_{\ell}\alpha_{\ell}W1_{\ell}-e^{-3i\pi/4}\frac{t^{3}}{6}\sum_{\ell}\alpha_{\ell}W2_{\ell}+O(t^{4}).$ (65)
By definition of $(p_{1},p_{2},p_{3})$, the polynomial $p(t)$ is the negative of the displayed cubic truncation, so
$A(u(t))+p(t)=O(t^{4}),\qquad t\to 0\ \text{on}\ \Omega_{+}^{\mathrm{phys}}.$ (66)
The identical argument with $\beta_{\ell}$ yields
$B(u(t))+q(t)=O(t^{4}).$ (67)
9 $\tau^{2}$ pairing compression
Define the involution on labels
$\ell=(m,\sigma,\varepsilon_{w})\longmapsto\ell^{\prime}=(m+2\bmod 4,\sigma,\varepsilon_{w}).$ (68)
Under this pairing, the pole transport implies
$t_{\ell^{\prime}}=-t_{\ell},\qquad Y_{\ell^{\prime}}=Y_{\ell},$ (69)
and the phase symbols flip:
$\chi_{m+2}=-\chi_{m},\qquad\psi_{m+2}=-\psi_{m},\qquad\kappa_{m+2}=-\kappa_{m}.$ (70)
Applying the derivative convention (30) under $t\mapsto-t$ yields $w^{\prime}_{m+2}(u_{\ell^{\prime}})=-w^{\prime}_{m}(u_{\ell})$ and hence
$r_{I}(\ell^{\prime})=-r_{I}(\ell).$ (71)
Using the half-period identities (42), we have $W0_{\ell^{\prime}}=W0_{\ell}$, $W2_{\ell^{\prime}}=W2_{\ell}$, but $W1_{\ell^{\prime}}=-W1_{\ell}$. Moreover, from the coefficient tables in Proposition 6.16.2, the phase flip and $r_{I}$ flip cancel, so $\alpha_{\ell^{\prime}}=\alpha_{\ell}$ and $\beta_{\ell^{\prime}}=\beta_{\ell}$.
Consequently:
  • Pair contributions double in $p_{1},p_{3},q_{1},q_{3}$ (built from $W0$ and $W2$).
  • Pair contributions cancel in $p_{2},q_{2}$ (built from $W1$).
The only broken pair arises from excluding $\ell_{\mathrm{inc}}=(0,+,-)$, whose partner is $\ell^{\prime}_{\mathrm{inc}}=(2,+,-)$, so in particular
$p_{2}=p_{2}^{(\ell^{\prime}_{\mathrm{inc}})},\qquad q_{2}=q_{2}^{(\ell^{\prime}_{\mathrm{inc}})}.$ (72)
10 Singular channel, explicit $d_{\ell}$ table, and canonical no-double-counting theorem
10.1 Singular channel and residue definition
Define
$\beta_{\mathrm{ch}}(t,Y):=\frac{iY}{\sqrt{2}}(t^{2}-1),$ (73)
and the singular channel
$P_{1,3}(u)=\frac{A(u)+p(t(u))}{4t(u)^{4}}-\beta_{\mathrm{ch}}(t(u),Y(u))\frac{B(u)+q(t(u))}{4t(u)^{4}}.$ (74)
At each forcing pole $u=u_{\ell}$, we have $\operatorname*{Res}_{u=u_{\ell}}A(u)=\alpha_{\ell}$ and $\operatorname*{Res}_{u=u_{\ell}}B(u)=\beta_{\ell}$, hence the residue
$d_{\ell}:=\operatorname*{Res}_{u=u_{\ell}}P_{1,3}(u)=\frac{\alpha_{\ell}-\beta_{\mathrm{ch}}(t_{\ell},Y_{\ell})\beta_{\ell}}{4t_{\ell}^{4}}.$ (75)
10.2 Explicit $d_{\ell}$ parity $\times j$ table
Substitute Proposition 6.16.2 into (75) and simplify using only $Y^{2}=2(t^{4}+1)$.
Proposition 10.1 (Explicit $d_{\ell}$ table).
Write $(t,Y)=(t_{\ell},Y_{\ell})$ and $r_{I}=r_{I}(\ell)$.
$\text{If $m$ is even:}\quad d_{\ell}$ $=\begin{cases}\displaystyle\frac{\chi_{m}r_{I}}{4t^{2}}\bigl(t^{6}-t^{4}+t^{2}-2\bigr),&j=1,\\[5.16663pt] \displaystyle\frac{\chi_{m}r_{I}}{4t^{4}}\bigl(t^{6}-2t^{4}+2t^{2}-2\bigr),&j=2,\\[5.16663pt] \displaystyle\frac{\chi_{m}r_{I}}{4}\,\frac{iY}{\sqrt{2}}(t^{2}-1),&j=3,\\[5.16663pt] 0,&j=4;\end{cases}$
$\text{If $m$ is odd:}\quad d_{\ell}$ $=\begin{cases}\displaystyle\frac{\psi_{m}r_{I}}{4},&j=1,\\[4.30554pt] \displaystyle\frac{\psi_{m}r_{I}}{4}\,t^{4},&j=2,\\[4.30554pt] \displaystyle\frac{\kappa_{m}r_{I}}{4}\,\frac{Y}{\sqrt{2}}\,\frac{2-t^{2}}{t^{2}},&j=3,\\[5.16663pt] \displaystyle\frac{\kappa_{m}r_{I}}{4}\,\frac{Y}{\sqrt{2}}\,\frac{t^{6}-2}{t^{4}},&j=4.\end{cases}$
Remark 10.2 (Flagged mechanism).
In the cases (meven,j=1)(m\ $\text${even},j=1), (meven,j=2)(m\ $\text${even},j=2), and (modd,j=2)(m\ $\text${odd},j=2), the $\beta_{\ell}$ carries a factor $Y$, so $\beta_{\mathrm{ch}}(t,Y)\beta_{\ell}$ carries $Y^{2}$. Replacing $Y^{2}$ by $2(t^{4}+1)$ via $Y^{2}=2(t^{4}+1)$ removes $Y$, making these $d_{\ell}$ purely $t$-rational, as visible in Proposition 10.1.
10.3 Remainder $R(u)$ and pole cancellation
Define
$R(u):=P_{1,3}(u)-\sum_{\ell\in I_{\mathrm{scat}}}d_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr].$ (76)
Lemma 10.3 (Pole cancellation and analyticity of $R$).
(i) $R(u)$ has no pole at any $u=u_{\ell}$, $\ell\in I_{\mathrm{scat}}$.
(ii) $R(u)$ is analytic at $u=u_{0}$ (equivalently at $t=0$ on $\Omega_{+}^{\mathrm{phys}}$).
Proof.
(i) By definition (75), the principal part of $P_{1,3}$ at $u_{\ell}$ is $d_{\ell}/(u-u_{\ell})$. The zeta difference $\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})$ has principal part $1/(u-u_{\ell})$. Subtracting $d_{\ell}[\cdots]$ cancels the principal part at each $u_{\ell}$.
(ii) By jet cancellation, $A(u(t))+p(t)=O(t^{4})$ and $B(u(t))+q(t)=O(t^{4})$ as $t\to 0$, so each fraction in $P_{1,3}$ is analytic at $t=0$. Each zeta difference vanishes at $u=u_{0}$, hence is analytic there. Therefore $R$ is analytic at $u_{0}$. ∎
10.4 Canonical no-double-counting representation
Theorem 10.4 (Canonical decomposition of $Q_{\mathrm{scat}}$).
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4. Let the coefficient tables $(\alpha_{\ell})$, $(\beta_{\ell})$ and $(C_{\ell})$ be given by Propositions 6.16.3, and define $(d_{\ell})$ by Proposition 10.1.
$Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr]+R(u),$ (77)
where $C_{\ell}$ is given by Proposition 6.3 and $R$ is as above. Then $Q_{\mathrm{scat}}$ has poles exactly at the forcing points $u_{\ell}$ ($\ell\in I_{\mathrm{scat}}$) with residues $C_{\ell}$, and the remainder $R$ is pole-free at every $u_{\ell}$ and analytic at $u_{0}$. In particular, the decomposition contains no double counting of principal parts from the singular channel $P_{1,3}$.
Theorem 10.5 (Uniqueness in the Sommerfeld class).
Assume the analytic and growth hypotheses on Sommerfeld densities from Lemma 1.3, together with the gauge normalization $Q_{\mathrm{scat}}(u_{0})=0$. Let $Q_{\mathrm{scat}}$ be the scattered spectral density constructed in Theorem 10.4. If $\widetilde{Q}_{\mathrm{scat}}$ is any other meromorphic function on the physical branch with at most simple poles at the forcing points $\{u_{\ell}:\ \ell\in I_{\mathrm{scat}}\}$, satisfying the same spectral functional system and scattered allocation, and the same gauge normalization $\widetilde{Q}_{\mathrm{scat}}(u_{0})=0$, then $\widetilde{Q}_{\mathrm{scat}}\equiv Q_{\mathrm{scat}}$. Equivalently, within this Sommerfeld class the corresponding scattered fields are unique.
Proof.
Let $H:=\widetilde{Q}_{\mathrm{scat}}-Q_{\mathrm{scat}}$. By linearity of the functional system, $H$ satisfies the associated homogeneous system (zero forcing), and by the local residue relations in Appendix A its principal parts at each forcing point are uniquely determined. Since both $\widetilde{Q}_{\mathrm{scat}}$ and $Q_{\mathrm{scat}}$ satisfy the same residue tables (Propositions 6.16.3), the difference $H$ is analytic at every $u=u_{\ell}$ ($\ell\in I_{\mathrm{scat}}$). Moreover, $H$ has no jump across the contour system defining the additive Riemann–Hilbert problem, so the uniformization $u\mapsto(t,Y)$ implies that $H$ extends to a holomorphic elliptic function on the square torus. A holomorphic elliptic function is constant, hence $H\equiv c$. Finally, the gauge normalization gives $c=H(u_{0})=0$, so $H\equiv 0$. If the gauge is not imposed then $c$ is the only remaining ambiguity; this constant does not affect the Sommerfeld integrals because they involve the difference $Q(\theta+z)-Q(\theta-z)$ (cf. Lemma 1.3). ∎
11 Analyticity at the incident spectral point
Theorem 11.1 (Analyticity at the incident spectral point).
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4, and impose the scattered allocation by excluding the incident label $\ell_{\mathrm{inc}}=(0,+,-)$ from the inside set, i.e., $I_{\mathrm{scat}}=I\setminus\{\ell_{\mathrm{inc}}\}$. Then the scattered spectral density $Q_{\mathrm{scat}}(\zeta)$ is analytic at the incident spectral point $\zeta=\zeta_{i}=\theta_{i}+i\varepsilon$ (for each fixed $\varepsilon>0$).
Proof.
By Theorem 10.4, the function $Q_{\mathrm{scat}}$ admits the canonical decomposition
$Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\bigr]+R(u),$
where each zeta difference has a simple pole only at $u=u_{\ell}$ and the remainder $R$ is pole-free at every forcing point. Since $\ell_{\mathrm{inc}}\notin I_{\mathrm{scat}}$, no term in the zeta sum has a pole at the incident point $u=u_{\ell_{\mathrm{inc}}}$. Moreover, the definitions of $A(u)$, $B(u)$ and hence of $P_{1,3}(u)$ and $R(u)$ involve sums only over $I_{\mathrm{scat}}$, so $R(u)$ is analytic at $u=u_{\ell_{\mathrm{inc}}}$ as well. Therefore $Q_{\mathrm{scat}}$ is analytic at $u=u_{\ell_{\mathrm{inc}}}$, and hence, by the physical lift $u=u(\zeta)$, analytic at $\zeta=\zeta_{i}$. ∎
12 Radiation condition and Meixner edge condition in the spectral formulation
The Sommerfeld representations (2)–(3) are classical in wedge diffraction. For completeness we record sufficient hypotheses on the spectral densities $Q$ and $S$ ensuring (i) the Sommerfeld radiation condition as $r\to\infty$ and (ii) the Meixner finite-energy edge condition at the wedge apex $r\to 0$. Statements of this type are standard; see, for example, Noble [4, Chs. 2–4] and Rawlins [6, § 3] (and the original half-plane analysis of Sommerfeld [1]).
Proposition 12.1 (Radiation and Meixner conditions from Sommerfeld data).
Assume that there exists $\eta>0$ such that the densities $Q(\zeta)$ and $S(\zeta)$ are meromorphic in the strip
$\mathcal{S}_{\eta}:=\{\zeta\in\mathbb{C}:|\Im\zeta|<\eta\},$ (78)
with at most finitely many simple poles, all displaced away from the integration contour $\gamma$ by the limiting-absorption prescription $\varepsilon>0$. Assume also that for some constants $C,N$ one has the uniform growth bound
$|Q(\zeta)|+|S(\zeta)|\leq C(1+|\zeta|)^{N},\qquad\zeta\in\mathcal{S}_{\eta}\setminus\{\text{poles}\}.$ (79)
Finally, assume a gauge normalization on the physical branch, for example
$Q_{\mathrm{scat}}(u_{0})=0,\qquad\text{equivalently }\quad Q_{\mathrm{scat}}(\zeta)\to 0\ \text{as }\Im\zeta\to+\infty,$ (80)
(and likewise for $S_{\mathrm{scat}}$).
Then the Sommerfeld integrals (2)–(3) define classical solutions of the Helmholtz equations in their respective sectors, satisfy the Sommerfeld radiation condition as $r\to\infty$, and satisfy the Meixner finite-energy condition at the wedge apex $r\to 0$.
Proof.
Under (78)–(79) the contour $\gamma$ can be deformed within the strip to the standard pair of rays with $\Im z>0$ and $\Im z<0$ without crossing singularities (cf. [4, Ch. 2]). The resulting integrals converge absolutely and allow differentiation under the integral sign; hence the reconstructed fields solve the Helmholtz equations in each sector.
For $r\to\infty$, steepest descent on the phase $\cos z$ along the deformed contour yields an outgoing leading term proportional to $e^{ikr}/\sqrt{r}$, with remainder $o(r^{-1/2})$; see, for example, [17, Ch. 6] or [18, § 2.4] for general stationary-phase estimates, and [3] for the diffraction interpretation in wedge geometries. The outgoing far-field expansion implies the Sommerfeld radiation condition.
For $r\to 0$, one expands $e^{ikr\cos z}=1+O(r)$ uniformly on $\gamma$ and uses (80) together with the strip analyticity to shift the contour upward, obtaining boundedness of $u$ and its first derivatives in a neighborhood of the apex; boundedness of $u$ and $\nabla u$ implies the Meixner finite-energy condition (see [6, § 3] and the discussion in [4, Ch. 3]). ∎
Corollary 12.2 (Radiation and Meixner for the constructed solution).
The densities $Q_{\mathrm{scat}}$ and $S_{\mathrm{scat}}$ constructed in §10 satisfy the hypotheses of Proposition 12.1. Consequently, the fields reconstructed by (2)–(3) satisfy the Sommerfeld radiation condition and the Meixner edge condition.
Proof.
[…]
The only singularities of $Q_{\mathrm{scat}}$ and $S_{\mathrm{scat}}$ in the strip are the prescribed simple poles (with $\varepsilon>0$ displacing them away from $\gamma$), and $Q_{\mathrm{scat}}$, $S_{\mathrm{scat}}$ are $2\pi$-periodic in $\Re\zeta$ away from poles because $s_{\zeta}=e^{i(\zeta-\theta_{w})}$ is $2\pi$-periodic. Hence the growth condition (79) holds (in fact with $N=0$) on compact subsets of the strip avoiding the poles. Proposition 12.1 applies. ∎
13 Far-field diffraction coefficient
Steepest-descent justification.
The contour deformation and stationary-phase evaluation used in Section 13 are justified under standard analyticity and growth hypotheses for $Q_{\mathrm{scat}}(\zeta)$ in a strip containing the real axis; see, for example, Bleistein–Handelsman [17, §2.4] or Wong [18, Ch. II]. In the present lemniscatic case these hypotheses are met for $\varepsilon>0$ because $Q_{\mathrm{scat}}$ is given by an explicit elliptic-function representation (Theorem 10.4) and the forcing poles are displaced off the real $\zeta$–axis by the limiting absorption parameter.
Theorem 13.1 (Diffraction coefficient).
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4. Let the coefficient tables $(\alpha_{\ell})$, $(\beta_{\ell})$ and $(C_{\ell})$ be given by Propositions 6.16.3, and define $(d_{\ell})$ by Proposition 10.1.
$\boxed{\ D(\theta,\theta_{i})=e^{-i3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\ Q_{\mathrm{scat}}(\theta)\ }.$ (81)
Here $Q_{\mathrm{scat}}(\theta)$ denotes the physical boundary value of $Q_{\mathrm{scat}}(\zeta)$ at $\zeta=\theta$, obtained by evaluating the physical lift $u(\zeta)$ and taking the limiting absorption limit $\varepsilon\to 0^{+}$ at the end.
Proof.
Start from the Sommerfeld representation (2) for the scattered field in medium 0,
$u_{\mathrm{scat}}^{(0)}(r,\theta)=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{0}r\cos z}\Bigl(Q_{\mathrm{scat}}(\theta+z)-Q_{\mathrm{scat}}(\theta-z)\Bigr)\,dz.$ (82)
In the second term substitute $z\mapsto-z$ (so $\cos z$ is unchanged) to obtain
$u_{\mathrm{scat}}^{(0)}(r,\theta)=\frac{1}{2\pi i}\left(\int_{\gamma}+\int_{-\gamma}\right)e^{ik_{0}r\cos z}\,Q_{\mathrm{scat}}(\theta+z)\,dz.$ (83)
Under the analyticity and strip-growth hypotheses stated above, the union $\gamma\cup(-\gamma)$ may be deformed to the steepest descent rays through the saddle $z=0$, where $\cos z=1-\tfrac{1}{2}z^{2}+O(z^{4})$. The leading contribution is therefore
$u_{\mathrm{scat}}^{(0)}(r,\theta)\sim\frac{2}{2\pi i}\,Q_{\mathrm{scat}}(\theta)\,e^{ik_{0}r}\int_{-\infty}^{\infty}\exp\!\left(-i\frac{k_{0}r}{2}x^{2}\right)\,\mathrm{d}x,\qquad r\to\infty.$ (84)
Using $\int_{-\infty}^{\infty}e^{-iax^{2}}\,\mathrm{d}x=\sqrt{\pi/a}\,e^{-i\pi/4}$ for $a>0$ yields
$u_{\mathrm{scat}}^{(0)}(r,\theta)\sim e^{ik_{0}r}\,\frac{D(\theta,\theta_{i})}{\sqrt{r}},\qquad D(\theta,\theta_{i})=e^{-i3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$ (85)
which is the claimed diffraction coefficient. ∎
This normalization is consistent with the standard two-dimensional GTD convention for wedge diffraction [3].
14 Conclusion and outlook
This paper provides an explicit elliptic-function reconstruction for the Sommerfeld spectral density $Q_{\mathrm{scat}}$ in the impedance-matched right-angle penetrable wedge at refractive index ratio $\nu=\sqrt{2}$. The construction is algebraic on the lemniscatic Snell surface $\Sigma_{\mathrm{lem}}$ and is written in terms of finite Weierstrass zeta differences on the square lattice together with an explicitly constructed holomorphic remainder that removes the partial-index singular channel without double counting.
What is proved/constructed.
  • A Sommerfeld spectral representation for the transmission problem is reduced to a closed two-face functional system in the spectral variable (Section 1.1).
  • In the integrable configuration $(\theta_{w},\nu,\rho)=(\pi/4,\sqrt{2},1)$, the spectral map closes on the lemniscatic curve $\Sigma_{\mathrm{lem}}$ and is uniformized by square-lattice Weierstrass functions (Sections 23).
  • The jet-killing polynomials $p,q$ are constructed so that $A(u)+p(t(u))=\mathcal{O}(t(u)^{4})$ and $B(u)+q(t(u))=\mathcal{O}(t(u)^{4})$ on $\Omega_{+}^{\mathrm{phys}}$, yielding analyticity at the physical basepoint $u_{0}$ (Section 8).
  • A canonical no-double-counting decomposition $Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})]+R(u)$ is obtained, with $R$ pole-free at all forcing points and analytic at $u_{0}$ (Theorem 10.4).
  • The far-field diffraction coefficient is expressed in terms of the physical boundary value $Q_{\mathrm{scat}}(\theta)$ (Theorem 13.1).
Limitations.
  • The result is restricted to the integrable lemniscatic regime $(\theta_{w},\nu,\rho)=(\pi/4,\sqrt{2},1)$; it does not address general wedge angles, general contrast, or non-impedance-matched media.
  • Outside special closures of the Snell surface, the spectral reductions typically lead to matrix Wiener–Hopf/Riemann–Hilbert factorization problems that are not treated here.
  • The present explicit tables are derived for the right-angle configuration; their analogues for other parameters require new residue analysis.
Context and outlook.
Complete analytic solutions for penetrable (transmission) wedge diffraction are rare and, outside of special configurations, the spectral reductions typically lead to matrix or multi-variable factorization problems. Even the right-angled penetrable wedge has been treated primarily by semi-analytical and asymptotic methods; see Antipov–Silvestrov [10], Nethercote–Assier–Abrahams [12] and Kunz–Assier [11] for penetrable-wedge analyses, and Groth–Hewett–Langdon [13] for high-frequency numerical-asymptotic methods for penetrable convex polygons in which local corner diffraction is central. A natural direction is to identify other parameter regimes in which the Snell surface closes algebraically (possibly at higher genus) and to determine whether analogous jet-killing and residue-cancellation mechanisms can be carried out.
15 Symbolic evaluation recipe
Given $(\theta_{i},\varepsilon>0)$:
  1. 1.
    Enumerate all pole labels $\ell=(m,\sigma,\varepsilon_{w})$ with $m\in\{0,1,2,3\}$, $\sigma,\varepsilon_{w}\in\{\pm 1\}$, and remove $\ell_{\mathrm{inc}}=(0,+,-)$.
  2. 2.
    For each $\ell$, compute $j=j(\sigma,\varepsilon_{w})$ and $\varepsilon_{j}$.
  3. 3.
    For each $\ell$, compute $b=b_{m,\sigma,\varepsilon_{w}}$ from (24), then compute the inside root $t_{q}=t_{\mathrm{in}}(b)$ from (27) and set $Y_{q}=2bt_{q}-\sqrt{2}(t_{q}^{2}+1)$. Transport to $(t_{\ell},Y_{\ell})$ by (28).
  4. 4.
    Define $u_{\ell}$ (physical lift) by the uniformization $\wp(u_{\ell})=x_{W}(t_{\ell})$ and $\tfrac{1}{2}\wp^{\prime}(u_{\ell})=y_{W}(t_{\ell},Y_{\ell})$.
  5. 5.
    Compute $r_{I}(\ell)$ from (33).
  6. 6.
    Compute $\alpha_{\ell}$, $\beta_{\ell}$, and $C_{\ell}$ from Proposition 6.1, 6.2, and 6.3.
  7. 7.
    Compute $W0_{\ell}$, $W1_{\ell}$, $W2_{\ell}$ from (42).
  8. 8.
    Compute jet coefficients $(p_{1},p_{2},p_{3})$ and $(q_{1},q_{2},q_{3})$ from the defining sums, or equivalently by summing the canonical per-pole summands in Propositions 6.46.5 evaluated at $(t_{\ell},Y_{\ell})$.
  9. 9.
    Build $A(u)$ and $B(u)$ from their definitions.
  10. 10.
    Build $P_{1,3}(u)$ and compute $d_{\ell}$ from (75) or Proposition 10.1. Then form $R(u)$.
  11. 11.
    Evaluate $Q_{\mathrm{scat}}(u)$ via Theorem 10.4.
  12. 12.
    For a given $\zeta$, compute $(t(\zeta),Y(\zeta))$ from the physical branch of $s(t,Y)=s_{\zeta}$, lift to $u(\zeta)$ via the uniformization, and evaluate $Q_{\mathrm{scat}}(\zeta)=Q_{\mathrm{scat}}(u(\zeta))$.
  13. 13.
    Obtain the far-field diffraction coefficient from the diffraction theorem above.
Appendix A Derivation of the residue tables
This appendix explains how Propositions 6.16.3 are obtained from the global two-face spectral system in the lemniscatic configuration $(\theta_{w},\nu,\rho)=(\pi/4,\sqrt{2},1)$. The computation is finite: one reduces the two-face coupling relations to a four-point orbit system on the Snell surface, applies a length–$4$ discrete Fourier transform (DFT) to decouple the system into four $2\times 2$ mode problems, and then evaluates the forcing residues at each pole $u_{\ell}$.
A.1 Scope and provenance of the residue tables
The tables in Propositions 6.16.3 are not independent assumptions: they are explicit solutions of the residue-matching conditions obtained by taking residues of the mode system (86) at the forcing poles and propagating those residues through the inverse mode matrices $M_{U,k}^{-1}$. Once the mode matrices and the local coefficients $(A_{0},B_{0},A_{1},B_{1})$ are fixed, the derivation reduces to finite algebra.
All simplifications in this appendix use only:
  • the lemniscatic curve identity $Y^{2}=2(t^{4}+1)$,
  • the root-of-unity relations $\omega=i$ and $\omega^{4}=1$,
  • the definitions of $g^{\prime}(t,Y)$, $\tau(t,Y)$, and the local coefficients $(A_{0},B_{0},A_{1},B_{1})$.
In particular, we do not invoke the additional pole relations $Y=2bt-\sqrt{2}(t^{2}+1)$ used in constructing the poles themselves; the residue tables are identities on the Snell surface.
A.2 Orbit reduction and mode matrices
For the right-angle wedge the two faces differ by a rotation of $2\theta_{w}=\pi/2$. On the lemniscatic surface this rotation is implemented by the automorphism $\tau(t,Y)=(it,-Y)$, see §2. After orbit closure one obtains, for each DFT mode $k\in\{0,1,2,3\}$, a matrix Wiener–Hopf/Riemann–Hilbert jump relation of the form
$M_{U,k}(t,Y)\,\widehat{U}_{k}^{+}(t,Y)=M_{V,k}(t,Y)\,\widehat{U}_{k}^{-}(t,Y)+\widehat{H}_{k}(t,Y),\qquad(t,Y)\in\Gamma,$ (86)
where $\Gamma=\{|t|=1\}$ is the physical cut, $\widehat{U}_{k}^{\pm}$ denote the boundary values on the two sides of $\Gamma$, and $\widehat{H}_{k}$ is the DFT forcing term generated by the incident wave. A derivation of such an orbit/DFT reduction for penetrable wedge systems is standard; see, for example, [6].
In the impedance-matched case the face-coupling matrix in Proposition 1.4 depends only on the Snell derivative $w^{\prime}(z)=dw/dz$. In the lemniscatic uniformization one has $w^{\prime}(z)=g^{\prime}(t,Y)$ with
$g^{\prime}(t,Y)=\sqrt{2}\,\frac{t^{2}-1}{Y},\qquad g^{\prime}(\tau(t,Y))=\sqrt{2}\,\frac{t^{2}+1}{Y},$
by (14) and $\tau(t,Y)=(it,-Y)$. Define
$A_{0}$ $:=\frac{1}{2}\bigl(1+g^{\prime}(t,Y)\bigr),$ $B_{0}$ $:=\frac{1}{2}\bigl(1-g^{\prime}(t,Y)\bigr),$
$A_{1}$ $:=\frac{1}{2}\bigl(1-g^{\prime}(\tau(t,Y))\bigr),$ $B_{1}$ $:=\frac{1}{2}\bigl(1+g^{\prime}(\tau(t,Y))\bigr),$
and let $\omega=i$. A convenient normalization of the mode coefficient matrices is
$M_{U,k}=\begin{pmatrix}-\omega^{-k}A_{1}&A_{0}\\[2.0pt] -\omega^{k}B_{1}&B_{0}\end{pmatrix},\qquad M_{V,k}=\begin{pmatrix}-\omega^{-k}B_{1}&B_{0}\\[2.0pt] -\omega^{k}A_{1}&A_{0}\end{pmatrix}$. (87)
$\Delta_{U,k}$ $=\det M_{U,k}=(-\omega^{-k}A_{1})B_{0}-A_{0}(-\omega^{k}B_{1})$ (88)
$=\omega^{k}A_{0}B_{1}-\omega^{-k}A_{1}B_{0},$
$M_{U,k}^{-1}$ $=\frac{1}{\Delta_{U,k}}\begin{pmatrix}B_{0}&-A_{0}\\[2.0pt] \omega^{k}B_{1}&-\omega^{-k}A_{1}\end{pmatrix}$.
$\Delta_{V,k}$ $=\det M_{V,k}=(-\omega^{-k}B_{1})A_{0}-B_{0}(-\omega^{k}A_{1})$ (89)
$=\omega^{k}A_{1}B_{0}-\omega^{-k}A_{0}B_{1},$
$M_{V,k}^{-1}$ $=\frac{1}{\Delta_{V,k}}\begin{pmatrix}A_{0}&-B_{0}\\[2.0pt] \omega^{k}A_{1}&-\omega^{-k}B_{1}\end{pmatrix}$.
A.3 Forcing residues and coefficient extraction
The forcing term $\widehat{H}_{k}$ is meromorphic on $\Sigma_{\mathrm{lem}}$ with simple poles at the forcing set $\{u_{\ell}\}$ defined in §4. Its residues are computed directly from the incident spectral density and the local phase $w_{m}(u)$ along the corresponding orbit branch. In particular, the scalar incident residue
$r_{I}(\ell)=\frac{\varepsilon_{j}}{w_{m}^{\prime}(u_{\ell})}$
is given explicitly by (33), and the phase factors $\chi_{m},\psi_{m},\kappa_{m}$ are as in (34).
For reproducibility, one may express the forcing residues in the mode variables in the following uniform form. Let $\omega=i$ and define $j=j(\sigma,\varepsilon_{w})$ by
$(+,+)\mapsto 3,\qquad(+,-)\mapsto 1,\qquad(-,+)\mapsto 4,\qquad(-,-)\mapsto 2.$
Let $(A_{m},B_{m})=(A_{0},B_{0})$ for $m$ even and $(A_{m},B_{m})=(A_{1},B_{1})$ for $m$ odd. Then
$\operatorname*{Res}_{u=u_{\ell}}\widehat{H}_{k}(u)=\omega^{-km}r_{I}(\ell)\,v_{k}^{(m,j)},$
where the 22–vector $v_{k}^{(m,j)}$ is given, for $j=1,2,3,4$, by
$\begin{array}[]{ll}j=1:&v_{k}^{(m,1)}=\bigl(\omega^{-k}A_{m},\;\omega^{k}B_{m}\bigr)^{\top},\\[2.84526pt] j=2:&v_{k}^{(m,2)}=\bigl(\omega^{-k}B_{m},\;\omega^{k}A_{m}\bigr)^{\top},\\[2.84526pt] j=3:&v_{k}^{(m,3)}=\bigl(-A_{m},\;-B_{m}\bigr)^{\top},\\[2.84526pt] j=4:&v_{k}^{(m,4)}=\bigl(-B_{m},\;-A_{m}\bigr)^{\top}.\end{array}$
Since the coefficient matrices $M_{U,k}$ are analytic and invertible at each forcing pole $u_{\ell}$ (the forcing poles occur away from the branch points), the jump relation (86) implies that the residue vector of the mode solution is obtained by solving a $2\times 2$ linear system:
$g_{k}(\ell):=\operatorname*{Res}_{u=u_{\ell}}\widehat{U}_{k}(u)=M_{U,k}(u_{\ell})^{-1}\,\operatorname*{Res}_{u=u_{\ell}}\widehat{H}_{k}(u).$ (90)
The reconstruction coefficients used in the elliptic sum are extracted from these residue vectors. A convenient choice (matching the definitions in §6) is
$\alpha_{\ell}:=t_{\ell}^{4}\,e_{1}^{\top}g_{1}(\ell),\qquad\beta_{\ell}:=t_{\ell}^{4}\,e_{2}^{\top}g_{3}(\ell),\qquad C_{\ell}:=\frac{1}{4}\sum_{k=0}^{3}e_{1}^{\top}g_{k}(\ell),$ (91)
where $e_{1}=(1,0)^{\top}$ and $e_{2}=(0,1)^{\top}$.
Substituting (87)–(91) together with the explicit forcing residues above and simplifying using only the lemniscatic relation $Y^{2}=2(t^{4}+1)$ (and $\omega^{4}=1$) yields the closed forms for $\alpha_{\ell}$, $\beta_{\ell}$ and $C_{\ell}$ recorded in Propositions 6.16.3. No additional identities are used.
Lemma A.1 (Symbolic verification of the tables).
Fix a forcing label $\ell=(m,\sigma,\varepsilon_{w})$, form $j=j(\sigma,\varepsilon_{w})$, and evaluate (90)–(91) using the forcing residues and the explicit inverses (88). After reducing with $Y^{2}=2(t^{4}+1)$, the resulting expressions for $\alpha_{\ell}$, $\beta_{\ell}$ and $C_{\ell}$ coincide with the corresponding entries in Propositions 6.16.3.
Proof.
All quantities entering (90)–(91) are rational in $(t,Y)$ and $\omega$ once the lemniscatic curve constraint $Y^{2}=2(t^{4}+1)$ is imposed. For a fixed case $(m\bmod 2,j)$, insert the forcing residue vector $\operatorname*{Res}\widehat{H}_{k}(u_{\ell})$, compute $g_{k}(\ell)=M_{U,k}^{-1}(u_{\ell})\operatorname*{Res}\widehat{H}_{k}(u_{\ell})$, and then read off $t_{\ell}^{4}e_{1}^{\top}g_{1}(\ell)$, $t_{\ell}^{4}e_{2}^{\top}g_{3}(\ell)$ and $\frac{1}{4}\sum_{k=0}^{3}e_{1}^{\top}g_{k}(\ell)$. The subsequent simplification is algebraic and uses only $Y^{2}=2(t^{4}+1)$ and $\omega^{4}=1$. ∎
A.4 Worked examples
We illustrate the residue extraction in two representative cases. Throughout we use only the matrix definitions (87)–(88), the forcing-residue vectors described above, and the lemniscatic identity $Y^{2}=2(t^{4}+1)$.
Example A (even $m$, case $m=0$, $j=1$).
Here $\chi_{0}=1$. For $k=1$ and $m=0$ the forcing residue vector (case $j=1$) is
$\operatorname*{Res}\widehat{H}_{1}(u_{\ell})=r_{I}(\ell)\begin{pmatrix}\omega^{-1}A_{0}\\[1.0pt] \omega B_{0}\end{pmatrix}=r_{I}(\ell)\begin{pmatrix}-iA_{0}\\[1.0pt] iB_{0}\end{pmatrix}$. (92)
Using (88) we obtain
$e_{1}^{\top}g_{1}(\ell)=\frac{1}{\Delta_{U,1}}\bigl(B_{0}(-iA_{0})+(-A_{0})(iB_{0})\bigr)r_{I}(\ell)=\frac{-2iA_{0}B_{0}}{\Delta_{U,1}}\,r_{I}(\ell).$ (93)
Since $\omega=i$ and $\omega^{-1}=-i$, $\Delta_{U,1}=\omega A_{0}B_{1}-\omega^{-1}A_{1}B_{0}=i(A_{0}B_{1}+A_{1}B_{0})$. Substituting the explicit lemniscatic coefficients $A_{0}=\tfrac{1}{2}(1+\sqrt{2}(t^{2}-1)/Y)$, $B_{0}=\tfrac{1}{2}(1-\sqrt{2}(t^{2}-1)/Y)$, $A_{1}=\tfrac{1}{2}(1-\sqrt{2}(t^{2}+1)/Y)$, $B_{1}=\tfrac{1}{2}(1+\sqrt{2}(t^{2}+1)/Y)$ into (93) and reducing with $Y^{2}=2(t^{4}+1)$ yields
$e_{1}^{\top}g_{1}(\ell)=-\frac{r_{I}(\ell)}{t_{\ell}^{2}},\qquad\alpha_{\ell}=t_{\ell}^{4}\,e_{1}^{\top}g_{1}(\ell)=-r_{I}(\ell)t_{\ell}^{2},$ (94)
which is the (meven,j=1)(m\ $\mathrm${even},j=1) entry in Proposition 6.1. A similar computation using (91) gives $C_{\ell}=-r_{I}(\ell)/(2t_{\ell}^{2})$, matching Proposition 6.3.
Example B (odd $m$, case $m=1$, $j=3$).
Here $\psi_{1}=i$ and $\kappa_{1}=-1$. For $k=1$ the forcing residue vector (case $j=3$) is
$\operatorname*{Res}\widehat{H}_{1}(u_{\ell})=\omega^{-1}r_{I}(\ell)\begin{pmatrix}-A_{1}\\[1.0pt] -B_{1}\end{pmatrix}=i\,r_{I}(\ell)\begin{pmatrix}A_{1}\\[1.0pt] B_{1}\end{pmatrix}$. (95)
Applying (88) and simplifying algebraically using $Y^{2}=2(t^{4}+1)$ gives
$e_{1}^{\top}g_{1}(\ell)=-\frac{Y_{\ell}}{\sqrt{2}\,t_{\ell}^{2}}\,r_{I}(\ell),\qquad\alpha_{\ell}=t_{\ell}^{4}\,e_{1}^{\top}g_{1}(\ell)=\kappa_{1}\,r_{I}(\ell)\,\frac{Y_{\ell}}{\sqrt{2}}\,t_{\ell}^{2},$ (96)
which coincides with the (modd,j=3)(m\ $\mathrm${odd},j=3) entry in Proposition 6.1. The remaining cases, including the $\beta_{\ell}$ and $C_{\ell}$ tables, follow from the same residue-insertion step combined with algebraic reduction on $\Sigma_{\mathrm{lem}}$.
A.5 Remark on symmetries
The case structure in Propositions 6.16.3 reflects the symmetries of the lemniscatic surface. The automorphism $\tau$ permutes orbit indices $m\mapsto m+1$ and exchanges the roles of the polynomials $(t^{2}-1)$ and $(t^{2}+1)$, while the involution $\eta(t,Y)=(t,-Y)$ flips the branch of $Y$. These symmetries reduce the residue computation to a small number of representative configurations.